40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6043     Entry time: Tue Nov 29 19:08:53 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: EOM heater disconnected 

I disconnected the heater at ~2:20 UTC, leaving the sensor circuit operational. Don't be fooled by the apparent railing of the heater in the monitor trace below---the heater has been physically disconnected, so there is no current flowing even though the servo is railed (since the error signal is huge with the loop open).

Kiwamu and I also restarted c1sus and locked the MC so that we can get some uncontrolled Stochmon data. I think he is planning to reconnect the heater some hours from now so that we can get yet another controlled data stretch (since the first one was cut short by c1sus's going down).

Screen_shot_2011-11-29_at_7.03.36_PM.png

ELOG V3.1.3-